Evaluate f(x)=e^x for given x values:
(a) x=2, f(x)=
(b) x=-2, f(x)=
(c) x+3.5, f(x)=
(d) x=5/3, f(x)=
(Round to 4 decimals)

Answer by kkasko(46)

Evaluate f(x)=e^x for given x values:
press 2nd/ on the calculator to get e then press ^2=7.3891
(a) x=2, f(x)=7.3891
(b) x=-2, f(x)=.1353
(c) x+3.5, f(x)=x+3.5=0
x=-3.5
e^-3.5=.0302
(d) x=5/3, f(x)=5.2945
(Round to 4 decimals)